Range for the class 10-20, 20-30, 30-40? A) 10 B) 20 C) 30 D) 40​

Grade:12

2 Answers

Vikas TU
3 years ago
Dear student 
There is upper limit of class = 40 
Lower limit of class = 10 
So , Range of class = 40-10 = 30 
Option C is correct.,
Good Luck 
Cheers 
Sowmiya
2 years ago
Option C is correct There is upper limit of class = 40
Lower limit of class = 10
So , Range of class = 40-10 = 30
hence 
Option C is correct